OpenStudy (anonymous):

If x,y are the roots of ax^2+bx+c=0 , find the equation with roots (x-y)^2 , (x+y)^2 . . ANSWER= a^4x^2-2a^2(b^2-2ac)x+b^2(b^2-4ac)=0

hartnn (hartnn):

find x^2+y^2 then (x-y)^2 = x^2-2xy+y^2 you know xy ...and x+y so u know the roots.

OpenStudy (anonymous):

(x+y)^2-2xy-2xy+(x+y)^2-2xy+2xy

OpenStudy (anonymous):

is this rgight?

OpenStudy (anonymous):

*right

hartnn (hartnn):

what was that ?

OpenStudy (anonymous):

I expanded all

hartnn (hartnn):

x+y = -b/a xy = c/a x^2+y^2 =... ?

OpenStudy (anonymous):

x^2-2xy+y^2

hartnn (hartnn):

that is (x-y)^2 ...

OpenStudy (anonymous):

but we did it last time a^2+b^2=(a+b)^2-2ab

hartnn (hartnn):

ok... x^2+y^2 = (x+y)^2 -xy = b^2/a^2 - c/a got this ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

-xy?

hartnn (hartnn):

sorry, -2xy

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

b^2-2ac/a^2

hartnn (hartnn):

(x-y)^2 = x^2-2xy+y^2 = b^2/a^2 -2c/a -2c/a right ? and (x+y)^2 = b^2/a^2 so you now have the 2 roots!

OpenStudy (anonymous):

b^2-4ac/a^2

hartnn (hartnn):

and whats that ?

OpenStudy (anonymous):

mistake

OpenStudy (anonymous):

why 2c/a 2 times

OpenStudy (anonymous):

-2c/a-2c/a

hartnn (hartnn):

(x-y)^2 = x^2-2xy+y^2 =[x^2+y^2]-2[xy]= [ b^2/a^2 -2c/a] -2[c/a]

OpenStudy (anonymous):

(x+y)^2-2xy-2xy+(x+y)^2-2xy+2xy

hartnn (hartnn):

i am not sure, whats that^ you always only write an expression (one side) not equation

OpenStudy (anonymous):

i expanded all (x-y)^2 + (x+y)^2

hartnn (hartnn):

(x-y)^2 + (x+y)^2 = 2 (x^2+y^2)

OpenStudy (anonymous):

I think its wrong

OpenStudy (anonymous):

leave it... its wrong

OpenStudy (anonymous):

so we were here => b^2/a^2-2c/a-2c/a

OpenStudy (anonymous):

it forms b^2-2ac/a^2 am I correct?

hartnn (hartnn):

(b^2-4ac)/a^2

OpenStudy (anonymous):

yes exactly!! sorry

OpenStudy (anonymous):

what about (x+y)^2

hartnn (hartnn):

x+y= -b/a (x+y)^2 = b^2/a^2

OpenStudy (anonymous):

(b^2-4ac)/a^2 + b^2/a^2

OpenStudy (anonymous):

the denominator should be a^4

hartnn (hartnn):

for sum of roots, denominator = a^2 is correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the sum of roots is 2b^2-4ac / a^2

hartnn (hartnn):

yes.

OpenStudy (anonymous):

product of roots --> (x-y)^2 * (x+y)^2

OpenStudy (anonymous):

how to do it?

hartnn (hartnn):

you know both the roots... b^2/a^2 and (b^2-4ac)/a^2 just take this product...

OpenStudy (anonymous):

yes

hartnn (hartnn):

product of roots = b^2(b^2-4ac)/a^4 now can you get to the answer ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

ok, good :)